61) Assume that a company’s planned level of activity was 3,500 units and its actual level of activity was 4,000 units. The spending variance for one of its mixed expenses was $350 favorable and its activity variance was $200 unfavorable. The total amount of this mixed cost included in the planning budget was $12,000. What is the actual total amount of this mixed expense?

Multiple Choice

  • $12,250

  • $11,850

  • $11,450

  • $13,650

62)Assume that a company provided the following cost formulas for three of its expenses (where q refers to the number of hours worked):

Rent (fixed) $3,000
Supplies (variable) $4.00 q
Utilities (mixed) $150 + $0.75 q

The company’s planned level of activity was 2,100 hours and its actual level of activity was 1,850 hours. How much supplies expense would be included in the planning budget?
Multiple Choice

  • $7,800

  • $8,100

  • $8,300

  • $8,400

63) Assume that a company provided the following cost formulas for three of its expenses (where q refers to the number of hours worked):

Rent (fixed) $3,000
Supplies (variable) $4.80 q
Utilities (mixed) $150 + $0.75 q

The company’s planned level of activity was 2,000 hours and its actual level of activity was 1,850 hours. How much supplies expense would be included in the flexible budget?

Multiple Choice

  • $10,320

  • $8,880

  • $9,600

  • $9,180

64) Assume that a company provided the following cost formulas for three of its expenses (where q refers to the number of hours worked):

Rent (fixed) $3,000
Supplies (variable) $4.00 q
Utilities (mixed) $250 + $0.75 q

The company’s planned level of activity was 2,000 hours and its actual level of activity was 1,900 hours. How much utilities expense would be included in the planning budget?

Multiple Choice

  • $1,500

  • $1,675

  • $1,750

  • $1,775
Answer #1
61.) The actual total amount of this mixed expense $ 11,850 =12000-350+200
Correct answer is option 2.
62.) Supplies expense would be included in the planning budget $ 8,400 =2100*4
Correct answer is option 4.
63.) Supplies expense would be included in the flexible budget $ 8,880 =4.8*1850
Correct answer is option 2.
64.) Utilities expense would be included in the planning budget $ 1,750 =250+(0.75*2000)
Correct answer is option 3.